2. Determine the oxidation number for the indicated element in each of the following substances. a. So, b. Mno, c. HNO, d. Crthanks
Answer #1

For determining oxidation number of indicated elements , we will consider (-2) charge of oxygen atom and (+1) charge of hydrogen atom.

We will take 'x' for the indicated atom.

we know that charge of a neural molecule is zero.

a. 502 30 +.2001 - 0 + 2(-2) = 0 OC - 4-0 a = +4 b. mnou x + 4 (0) = -1 20 + 4 (-2)= -1 2C-8 :-1 OC :-) + 8 x = +7 C. HNO, d.

So Oxidation number for the indicated elements are:

a. S = +4

b.Mn = +7

c.N = +3

d.Cr = +6

Now,

4(a)

4.ca? we have, w= 0.75 gram v=850 ml molecular weight of Nu,soy m.w: 2 (na) + I (S) + 410) = 2(23)+1 (32) + 4(16) - 46 + 32 +

Molarity of a solution is 6.21× 10-3 M.

(b)

Cb) we have W: 2.5 gm (M): 1.5 m V: 9. molecular weight of NY,DOY m.w: 3 (NU) + ICP) + 410) = 3 (23) +1 (31) + 4(16) = 69 +31

Volume is 0.0102 liter.

PLEASE GIVE A THUMBS UP IF IT REALLY HELPS YOU.

THANK YOU.